40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 366     Entry time: Thu May 23 23:27:38 2019
Author: Koji 
Type: Optics 
Category: Characterization 
Subject: IGHQEX3000 high power test 

LaserComponents IGHQEX3000 (Cage B2: Serial# B1-23) was exposed to the beam with the optical power from 1.6mW to 332mW.
After each illumination, the dark current and the dark noise level were measured. Also the photo image of the PD surface was taken each time.

- No significant change of the dark current after each illumination.

- No significant change of the dark noise after each illumination.

- No visible change of the surface observed.

(During this dark noise measurement, the current amp gain was set to be 1e8 V/A, instead of 1e7 for the measurements yesterday.)

Attachment 1: HQEPD_high_power_test.pdf  183 kB  Uploaded Fri May 24 00:29:56 2019  | Hide | Hide all
HQEPD_high_power_test.pdf
Attachment 2: pd_images.png  3.565 MB  Uploaded Fri May 24 00:30:33 2019  | Hide | Hide all
pd_images.png
ELOG V3.1.3-